OpenStudy (anonymous):

The sum of the squares of two consecutive integers is 85. Find the integers.

hartnn (hartnn):

you know what consecutive integers are ?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

x^2+y^2=85. since you know they are consecutive, y must be 1 greater than x. so y=x+1. plug and play

hartnn (hartnn):

or initially you can take 2 numbers as x and x+1 and get x^2 + (x+1)^2 = 85 directly can you slve this quadratic ?

OpenStudy (anonymous):

then it would be 2x^2 + 2x - 84 = 0

hartnn (hartnn):

yes which is x^2+x-42 =0 can you factor ? 2 numbers with sum = 1 , product = -42 ?

OpenStudy (anonymous):

(x + 7) (x - 6) am i right?

hartnn (hartnn):

yes!

hartnn (hartnn):

i think you'll now get those 2 numbers :)

OpenStudy (anonymous):

so which will be the x then? will it be -7 or 6?

hartnn (hartnn):

both! -7, -7+1 -----> -7,-6 6, 6+1 ------> 6,7 :)

OpenStudy (anonymous):

okay. thanks :)

hartnn (hartnn):

welcome ^_^